Calculate the mass of sodium acetate (CH3COONa) required to make 500 mL of


0.375 molar aqueous solution. Molar mass of sodium acetate is 82.0245 g mol–1.


Grade:11

4 Answers

Radhika Batra
10 years ago

0.375 M aqueous solution of sodium acetate

≡ 1000 mL of solution containing 0.375 moles of sodium acetate

∴Number of moles of sodium acetate in 500 mL

 

Molar mass of sodium acetate = 82.0245 g mole–1 (Given)

∴ Required mass of sodium acetate = (82.0245 g mol–1) (0.1875 mole)

= 15.38 g

 

Lovish singla
5 years ago
Molarity°mass/volume; put molarity to be0.375 and volume to be 500 ml we will get moles to be .188 and then multiply by 82 mass of sodium acetate
aditya pradeep
5 years ago
0.375M aqueous solution means that 0.375 mo of sodium acetate are present in 1000mL of solution.500mL of the solution should contain sodium acetate.
                                                                         =0.375/2 mole
 molar mass of sodium acetate   =82.0245 gmole-1 
mass of sodium acetate required=(0.375/2) (82.0245)
                                                    =15.38 g 
 
ankit singh
3 years ago

0.375 M solution of CH3COONa = 0.375 moles of CH3COONa/1000ml

No. of moles of CH3COONa in 500ml =10000.375×500=0.1875 moles
Molar mass of sodium acetate = 82.0245 g/mole
mass of CH3COONa required =82.0245×0.1875=15.38g